# Call Me When Necessary: LLMs can Efficiently and Faithfully Reason over Structured Environments

### v1

@@ 5.2 Baselines
KGQA Baselines. Training-based methods fine-tune pre-tained language models (PLMs): EmbedKGQA Saxena et al. (2020), NSM He et al. (2021), TransferNet Shi et al. (2021), SR+NSM+E2E Zhang et al. (2022), UniKGQA Jiang et al. (2023c), ReasoningLM Jiang et al. (2023b) and RoG Luo et al. (2024). Inference-based methods call LLMs API for inference: Davinci-003 Ouyang et al. (2022), GPT3.5, GPT4 OpenAI (2023), AgentBench Liu et al. (2024b) and StructGPT Jiang et al. (2023a).
For fair comparison, we assume golden topic entities are given. Following Tan et al. (2023), we adopt Hit@1, assessing whether the predicted answer is correct, to evaluate KGQA.
TableQA Baselines. Training-based methods fine-tune PLMs: TAPAS Herzig et al. (2020), UnifiedSKG Xie et al. (2022), TAPEX Liu et al. (2022). Inference-based methods call LLMs API for inference: Davinci-003 Ouyang et al. (2022), GPT3.5, GPT4 OpenAI (2023), StructGPT Jiang et al. (2023a). Note that all Inference-based baselines regard the datasets as an information retrieval task.
We adopt denotation accuracy, assessing whether the predicted answer exactly matches the golden, to evaluate TableQA. Detailed descriptions of each baseline are in Appendix D.
@@ 5.3 Implementation Details
We adopt gpt-3.5-turbo OpenAI (2022) and GPT4 OpenAI (2023) as LLMs, denoted as Readi-GPT3.5 and Readi-GPT4. Temperature is 0.3 for all modules. For relation-binding, we deploy a Pyserini as a hybrid searcher with BM25 and Contriever Izacard et al. (2022). For each relation generated by LLMs, we retrieve top 5 similar relations on Freebase Bollacker et al. (2008). For instantiation, we deploy a Virtuoso server following the instructions [^2] https://github.com/dki-lab/Freebase-Setup. Prompt lists are in Appendix F.
@@ 5.4 Main Results
Results for KGQA As shown in Table 1, Overall, Readi significantly outperforms all inference-based methods, the vanilla LLMs and most training-based methods on all datasets, demonstrating the effectiveness of Readi. Compared with inference-based methods, Readi substantially boosts the results of the vanilla LLM (by 8.6% on WebQSP and 14.9% on CWQ), demonstrating that Readi enables LLMs to practically interact with structured environments. Moreover, Readi-GPT3.5 already significantly surpasses state-of-the-art results (by 4.7% on WebQSP and 12.4% on MQA-3H), and Readi-GPT4 further enhances the performance. With fewer LLM-calls, our directly-generated reasoning path and editing framework still achieves better performance. Compared with training-based methods, without large-scale supervision and the cost of beam search, Readi achieves comparable performance (e.g. 67.0 Hit@1 on CWQ) just by some demonstration examples, showing the faithfulness of Readi. Additionally, Readi sets new state-of-the-art results MQA-1H, illustrating the effectiveness of the editing module which refines the reasoning path if it goes wrong.
@@
Table 3: Answer Coverage Rate (AC) and QA Performance (Hit@1) of variance of Readi (GPT3.5 as reasoning module). Each row is by different initial path. Corrupt means corrupted path with some randomly-sampled relations. Empty means empty path. w/o edit means we remove the editing module and only leverage the initial reasoning path.
Results for TableQA We also experiment on TableQA scenario which requires multi-hop reasoning over table to show the generalizability of Readi. As shown in Table 2, overall, Readi outperforms all inference-based baselines and most training-based methods, setting state-of-the-art results on WTQ, demonstrating the effectiveness of our framework. Compared with inference-based methods, Readi surpasses previous state-of-the-art methods by 9.5% and the vanilla LLM by 5.9% on WTQ, demonstrating that LLMs perform well with directly generated reasoning path and some necessary editing. Compared with training-based methods, Readi is significantly superior on WTQ, while trailing behind on WikiSQL, without the cost of massive annotations. This may be due to the i.i.d. distribution between the training and testing sets of WikiSQL, which favors the results of fully-trained methods. Interestingly, Readi-GPT4 lags a little bit behind Readi-GPT3.5 on both datasets, which may be due to the information retrieval manner that asks LLMs to output answer directly based on retrieved table items, relying heavily on mathematical reasoning ability of LLMs, where GPT3.5 already plays on par with GPT4.
@@ 6 Analysis
We further analyze Readi’s modules, reasoning path and efficiency on 1000 test samples of CWQ. For fairness, we base all evaluation on our instantiation and reasoning modules.
@@ 6.1 Ablation Study
Effectiveness of reasoning path generation and editing modules. As shown in Table 3, we adpot the answer coverage rate (AC, rate of instances containing the answer) and QA performance (Hit@1) for illustration. We also test editing module with a corrupt path and an empty path. First, Readi establishes a plug-and-play nature for both modules, showing their effectiveness. With only an initial path (we instantiate the path and maintain the longest path if it goes wrong), Readi reaches comparable results (1st row). With an empty initial path (“Empty” columns), the editing modules still yield competitive results closer to the full Readi. Second, generally, higher capacity of LLMs leads to better results for both modules. This meets our expectation, suggesting future improvements of Readi. Third, Readi shows robustness for reasoning path editing, performing well even with an empty or even a corrupt path (“Corrupt” and “Empty” columns).
@@ 6.2 Reasoning Path Analysis
We compare reasoning path of Readi with representative fine-tuned methods, i.e., Subgraph Retrieval (SR) Zhang et al. (2022) trains an encoder to retrieve relations to build a path, RoG Luo et al. (2024) tunes a Llama 2 Touvron et al. (2023) to generate a path.
@@
   - beam size 1 58.4 26.3 50.9
   - beam size 3 67.2 47.1 54.6
@@
   - beam size 1 57.0 69.5 52.2
   - beam size 3 77.5 170.1 57.3
@@
How well Readi’s reasoning path is? We adopt answer coverage rate (AC) and number of retrieved knowledge (#RK) as the quality of graph. Ideally, the higher AC and lower #RK, the better. Also, we report Hit@1 as QA performance, shown in Table 4. First, Readi’s initial path is comparable with fine-tuned methods, with GPT4 surpassing them by a large margin (5.0% and 6.3% Hit@1 than RoG and SR, respectively), showing the effectiveness of our reasoning path. Second, with some necessary editing, Readi obtain substantially higher AC (while just a little higher #RK with GPT4 and even lower #RK with GPT3.5) and significantly higher QA performance than fine-tuned methods. Third, with wider beam size, fine-tuned methods gain higher AC, yet a drastically growing knowledge size, and lower Hit@1 than Readi, showing our superiority.
Extensive features of Readi’s reasoning path. The quality of reasoning path is multi-dimensional. To further show insights of reasoning path by LLMs, we design some metrics, including the instantiation progress, error types, etc. Also, we compare with SR Zhang et al. (2022), RoG Luo et al. (2024) and Golden (outermost in Figure 4). More detailed analysis are in Appendix E.
@@
Insights driven by Figure 4 are three-folded: 1) Readi’s initial path already achieves better in QA results. More than a half (60%) of Readi-GPT4’s initial paths can be fully instantiated (LLMs called only once), showing that Readi effectively unleash the intrinsic planning ability of LLMs. For each path, a large proportion (averagely 74%) can be instantiated, indicating the possibility to safe LLM-calls. 2) With necessary editing, Readi gets significantly closer to golden (from green to red in Figure 4), exceeding compared methods in many metrics, especially the QA results. 3) Interestingly, the path by LLMs establish a human-like nature, for humans tend to get stuck at points they have never seen in real world. For compound node ending rate (special for KG), fine-tuned methods are close to golden, showing that they memorize the structured information. However, they trail behind behind Readi on QA results, demonstrating that unfaithfulness still exits for fine-tuned methods.
@@ 7 Conclusion
In this paper, we propose a novel framework Readi where LLMs can reason over structured environments both efficiently and faithfully. In Readi, LLMs initially generate a reasoning path which is then instantiated on environments to facilitate faithful reasoning. Path editing is triggered only if the instantiation gets stuck. We showcase the implementation of Readi on knowledge graph, evaluate the effectiveness on KGQA and TableQA and analyze extensive features of Readi. Our work also shed lights on practically interaction between natural language and structured environments, where LLMs plays a crucial role to bridge the gap.
@@ 8 Limitations
Although Readi achieves strong performance in complex reasoning task over structured environments, e.g., knowledge graph and tables, there are still some limitations of our method. First, we leverage just 2 LLMs as backbone to evaluate Readi. Therefore, more experiments can be done to test performance of Readifor other LLMs (including the inference-based and training-based). Second, we adopt reasoning path as a representation of natural language utterance and model the task as a information-retrieval one. An interesting and more precise direction is semantic parsing, e.g., text2SQL, which we will leave for future work. Moreover, we can also test Readi on question answering over database or other reasoning fields. Third, the instantiation in our current implementation of KGQA is intuitive but may induce many queries for SPARQL engine if encountered some “big” entities. For example, when constraining “locations in France”, we may obtain hundreds of instances, for which we just set a threshold to restrict the number of current entities. This can be improved by meticulous query design. Forth, our proposal to fully harness intrinsic planning ability of LLMs performs well and saves some LLM-calls in experiments, which does not mean we should always depend on LLMs plans in every other task. And although the initial path by LLMs can be well instantiated, this just means the path by LLMs corresponds to some instances on KG, but not ensures the path exactly matches with the constraints in the question. Finally, by efficiency, we mean number of LLM-calls, or how LLMs can practically interact with structured environments. A promising direction is to reach for time and cost efficiency.
@@ Appendix A Implementation of Readi for TableQA
Lake Van Van Gölü 3755 km2 3755 km2 Van, Bitlis
Lake Tuz Tuz Gölü 1500 km2 2 m Aksaray, Ankara, Konya
Lake Palas Tuzla Palas Tuzla Gölü 106 km2 15 m Palas/Kayseri
Table 6: An example from WTQ Pasupat and Liang (2015), with the question “which is deeper, lake tuz or lake palas tuzla?”.
Algorithm 1 again demonstrates Readi’s framework. Our implementation of Readi for TableQA strictly sticks to Algorithm 1, and is just simpler than KGQA. The main difference between TableQA and KGQA is the reasoning path (mentioned in Section 3). Reasoning path for TableQA starts from a given table, to some columns and then to some rows, in order to constrain specific items in the table to answer the question. A sample of WTQ Pasupat and Liang (2015) dataset is shown in Table 6. The reasoning path of such question is from this table to column“Name in English” and “Depth”, and then to row items whose “Name in English” is “Lake Tuz” and “Lake Palas Tuzla”.
For reasoning path generation, we ask LLMs to generate a reasoning path for a given question and some table descriptions (including the header and a randomly-sampled example row), by in-context learning. Specifically, we require the LLMs to pick up at least to rows for complex arithmetic and aggregation operations in TableQA datasets.
@@ 9             Go to Line 11
For reasoning path instantiation, since schemas of a table is not as massive as a KG, we don’t need a relation binding module to retrieve candidate relations and connect them according to the reasoning path. For a table, instead, we first filter out columns and then filter out the rows in the reasoning path. If the selected columns or rows are not in the table, this is the "necessary" time to call LLMs to edit the path.
For reasoning path editing, similar to KGQA, we show the currently instantiated columns, and then give out candidate columns where the instantiation goes wrong, as error messages. Note that most questions of WTQ Pasupat and Liang (2015) and WikiSQL Zhong et al. (2017), so we just consider wrong columns as error. If a selected row in the reasoning path cannot be instantiated, we return all rows in the columns. Then we ask the LLMs to edit its reasoning path.
The QA reasoning is quite similar to KGQA, we concatenate the question and retrieved row items and ask the LLMs to answer the question.
@@ Appendix B Detailed Implementation of Readi for KGQA
The example prompts of reasoning path generation and reasoning path editing are in Appendix F. We demonstrate some details here.
@@ B.1 Reasoning Path Instantiation Details
In this part, we demonstrate the detailed path-connecting part mentioned in Section 4.3. Assume that we have a reasoning path starting from an entity, with $m$ relations $R$. All relations are binded to candidates $\hat{R}$. For the example in Figure 3, we have a reasoning path “[Nijmegen] serve_airport $\rightarrow$ contain” and have “$r_{1}:$ serve_airport” binded to “$\hat{r}_{1}:$ airport, terminal, serving port” in KG, and similarly have “$r_{2}:\leavevmode\nobreak$contain” binded to “$\hat{r}_{2}:\leavevmode\nobreak$contained by, contains, place of”. The purpose if this part is to check if there exists an instance in KG where the starting point is entity “Nijmegen”, then connect to some relations in $\hat{r}_{1}$, and then connect to some relations in $\hat{r}_{2}$.
The path-connecting is essentially an Breadth-first search (BFS) algorithm, with a queue $que$ containing possible entities connected to current relations. Firstly, we push the starting point (e.g., “Nijmegen”) into $que$. Second, at each time, we check if current candidate relations (e.g.,adjoin, terminal, serving port”) can connect to current head of $que$. If so, we successfully instantiate “serve_airport” in KG, and we can obtain some entities which are “serve_airport” of “Nijmegen” (e.g., “WZ air.” and “NTA.”). Then, we add these entities into the queue, and then match them with the rest relations.

### v2

@@ 5.2 Baselines
KGQA Baselines. Training-based methods fine-tune pre-trained language models (PLMs): EmbedKGQA Saxena et al. (2020), NSM He et al. (2021), TransferNet Shi et al. (2021), SR+NSM+E2E Zhang et al. (2022), UniKGQA Jiang et al. (2023c), ReasoningLM Jiang et al. (2023b) and RoG LUO et al. (2024). Inference-based methods call LLM-APIs: Davinci-003 Ouyang et al. (2022), GPT3.5, GPT4 OpenAI (2023), AgentBench Liu et al. (2024), StructGPT Jiang et al. (2023a). All baselines assume golden topic entities are given.
TableQA Baselines. Training-based methods find-tune PLMs: TAPAS Herzig et al. (2020), UnifiedSKG Xie et al. (2022), TAPEX Liu et al. (2022). Inference-based methods call LLM-APIs: Davinci-003 Ouyang et al. (2022), GPT3.5, GPT4 OpenAI (2023), StructGPT Jiang et al. (2023a). Note that Inference-based methods model TableQA as an information retrieval task.
Following Tan et al. (2023), we adopt Hit@1, assessing whether the predicted entity is correct, to evaluate KGQA. We adopt denotation accuracy, assessing whether the prediction exactly matches the golden, to evaluate TableQA. We elaborate baselines in Appendix C.3. We also discuss in-depth comparison with ToG Sun et al. (2024) and DATER Ye et al. (2023) in Appendix F.
@@ 5.3 Implementation Details
We adopt gpt-3.5-turbo OpenAI (2022) and GPT4 OpenAI (2023) as LLMs, denoted as Readi-GPT3.5 and Readi-GPT4. Temperature is 0.3 for all modules. For relation-binding, we deploy a Pyserini as a hybrid searcher with BM25 and Contriever Izacard et al. (2022). For each relation generated by LLMs, we retrieve top 5 similar relations on Freebase. For instantiation, we deploy a Virtuoso server following the instructions [^2] https://github.com/dki-lab/Freebase-Setup. More details can be found in Appendix C.
@@ 5.4 Main Results
Results for KGQA As shown in Table 1, overall, Readi significantly outperforms all Inference-based methods, the vanilla LLMs and most training-based methods on all datasets, demonstrating the effectiveness of Readi. Compared with inference-based methods, Readi substantially boosts the results of the vanilla LLM (by 8.6% on WebQSP and 14.9% on CWQ), demonstrating that Readi enables LLMs to practically interact with structured environments. Moreover, Readi-GPT3.5 already significantly surpasses state-of-the-art results with LLM-APIs (by 4.7% on WebQSP and 12.4% on MQA-3H), and Readi-GPT4 further enhances the performance. With fewer LLM-calls, our directly-generated reasoning path and editing framework still achieves better performance. Compared with Training-based methods, without large-scale supervision and cost of beam search, Readi achieves comparable performance (e.g. 67.0 Hit@1 on CWQ) by some demonstration examples, showing the effectiveness of Readi. Additionally, Readi sets new state-of-the-art results MQA-1H, showing the effectiveness of editing which provides pertinent feedback upon instantiation errors.
@@
Table 3: Answer Coverage Rate (AC) and QA Performance (Hit@1) of variance of Readi (GPT3.5 as reasoning module). Each column denotes a path generation method. Corrupt means a path with some randomly-sampled relations. Empty means empty path. w/o edit means we only leverage the initial reasoning path.
Results for TableQA We experiment on TableQA scenario requiring multi-hop reasoning over tables to show the generalizability of Readi. As shown in Table 2, overall, Readi outperforms all Inference-based and most Training-based baselines, setting state-of-the-art results on WTQ, demonstrating the effectiveness of our framework. Compared with Inference-based methods, Readi surpasses previous state-of-the-art methods by 9.5% and the vanilla LLM by 5.9% on WTQ, showing that Readi significantly improve the LLMs performance. Compared with Training-based methods, Readi, without massive annotations, is significantly superior on WTQ, while trailing behind on WikiSQL. This may be due to the i.i.d. distribution between the training and testing sets of WikiSQL, favoring the results of fully-trained methods. Interestingly, Readi-GPT3.5 are comparable with Readi-GPT4. This might because we asks LLMs to reason (with aggregation operations) directly based on retrieved table items. Further analysis shows that Readi-GPT3.5 has more chance of invoking editing than Readi-GPT4, which offers more pertinent feedback from the environments. Elaboration on effectiveness of editing is in Section 6.1, Section 6.2 and Appendix D.2.
@@ 6 Analysis
We further analyze Readi’s modules, reasoning path and efficiency on 1000 test samples of CWQ. For fairness, we base all evaluation on our instantiation and reasoning modules. Please refer to Appendix D for more detailed analysis of Readi.
@@ 6.1 Ablation Study
Effectiveness of reasoning path generation and editing modules. As shown in Table 3, we adpot the answer coverage rate (AC, rate of instances containing the answer) and QA performance (Hit@1) for illustration. We also analyze the robustness of editing module with a Corrupt path and an Empty path. First, Readi establishes a plug-and-play nature for both modules, showing their effectiveness. With only an initial path (we instantiate the path and maintain the longest path if it goes wrong), Readi reaches comparable results (1st row). With an Empty initial path (Empty columns), the editing modules still yield competitive results close to the full Readi. Second, editing with LLMs significantly improves the performance (2nd and 3rd rows), further showing the effectiveness of the editing module. Third, generally, higher capacity of LLMs leads to better results for both modules, which meets our expectation. Lastly, Readi shows robustness for reasoning path editing, performing well even with an Empty or even a Corrupt path (Corrupt and Empty columns).
@@ 6.2 Reasoning Path Analysis
We compare reasoning path of Readi with representative fine-tuned methods, i.e., Subgraph Retrieval (SR) Zhang et al. (2022) trains an encoder to retrieve relation and then builds a path, RoG LUO et al. (2024) tunes a Llama 2 Touvron et al. (2023) to generate a path.
@@
- beam size 1 58.4 26.3 50.9
- beam size 3 67.2 47.1 54.6
@@
- beam size 1 57.0 69.5 52.2
- beam size 3 77.5 170.1 57.3
@@
Quality of Readi’s reasoning path. We adopt answer coverage rate (AC) and number of retrieved knowledge (#RK) as the quality of graph. Ideally, the higher AC and lower #RK, the better. Also, we analyze the QA performance (Hit@1), shown in Table 4. First, Readi’s initial path is comparable with fine-tuned methods, with GPT4 surpassing them by a large margin (5.0% and 6.3% Hit@1 than RoG and SR, respectively), showing the effectiveness of our reasoning path. Second, with some necessary editing, Readi obtains substantially higher AC, with a little higher #RK (GPT4) and even lower #RK (GPT3.5), and significantly higher QA performance than fine-tuned methods, showing the effectiveness of LLMs editing. Third, with wider beam size, fine-tuned methods gain higher AC, yet a drastically growing #RK, and lower Hit@1 than Readi, illustrating our superiority.
Extensive features of Readi’s reasoning path. The quality of reasoning path is multi-dimensional. To further show insights of reasoning path by LLMs, we design some metrics, including the instantiation progress, error types, etc. We compare with SR, RoG and Golden (Outermost in Figure 4). More detailed analysis is in Appendix D.1.
@@
Insights driven by Figure 4 are three-folded: 1) Readi’s initial path already achieves better in QA results. More than a half (60%) of Readi’s initial paths can be fully instantiated (LLMs called only once), showing that Readi effectively unleash the strong planning ability of LLMs. For each path, a large proportion (averagely 74%) can be instantiated, again indicating the possibility to safe LLM-calls. 2) With necessary editing, Readi gets significantly closer to golden (From green to red in Figure 4), exceeding compared methods in many metrics, especially the QA results. 3) Interestingly, the path by LLMs establish a human-like nature, for humans tend to get stuck at points they have never seen in real world. For Compound Node Ending Rate (special for KG), fine-tuned methods are close to Golden, showing that they memorize the structured information. However, they trail behind behind Readi on QA results, demonstrating that unfaithfulness still exits for fine-tuned methods.
@@ 7 Conclusion
In this paper, we propose a novel framework Readi where LLMs can reason over structured environments efficiently and faithfully. In Readi, LLMs initially generate a reasoning path which is then instantiated on environments to facilitate faithful reasoning. Path editing is triggered only if the instantiation gets stuck. We showcase the implementation of Readi on knowledge graph, evaluate the effectiveness on KGQA and TableQA and analyze extensive features of Readi. Our work also shed lights on practically interaction between natural language and structured environments, where LLMs plays a crucial role to bridge the gap.
@@ 8 Limitations
Although Readi achieves strong performance in complex reasoning task over structured environments, e.g., knowledge graph and tables, there are still some limitations of our method. First, we leverage only two LLMs as backbone to evaluate Readi. Therefore, more experiments can be done to test performance of Readi for other LLMs (including the inference-based and training-based). Second, we adopt reasoning path as a representation of natural language utterance and model the task as a information-retrieval one. An interesting direction is semantic parsing, e.g., text2SQL, which we will leave for future work. Moreover, we can also test Readi on question answering over database or other reasoning fields. Third, the instantiation in our current implementation of KGQA is intuitive but may induce many queries for SPARQL engine if encountered some “big” entities. Finally, our proposal, to fully harness intrinsic planning ability of LLMs and incorporate structural reasoning log to edit the path, performs well and saves some LLM-calls in experiments, which does not mean we should always depend on LLMs plans. And although the initial path by LLMs can be well instantiated, this just means the path by LLMs corresponds to some instances on KG, but not ensures that the path exactly matches with the constraints in the question.
9 Acknowledgements
The authors would like to thank Jue Zhang, Yuzhong Qu, Dapeng Li, Jinmao Li, Kaikai An, Yurong Wu and Wentao Ding for their valuable feedback. The authors would also like to thank Xiaokang Zhang for providing the reasoning path of Subgraph Retrieval.
@@ Appendix A Implementation of Readi for TableQA
Lake Van Van Gölü 3755 $km^{2}$ 171 m Van, Bitlis
Lake Tuz Tuz Gölü 1500 $km^{2}$ 2 m Aksaray, Ankara, Konya
Lake Palas Tuzla Palas Tuzla Gölü 106 $km^{2}$ 15 m Palas/Kayseri
Table 6: An example from WTQ Pasupat and Liang (2015), with the question ‘‘which is deeper, lake tuz or lake palas tuzla?’’.
Algorithm 1 demonstrates Readi’s framework. Our implementation of Readi for TableQA strictly sticks to Algorithm 1, and is simpler than KGQA. The main difference between TableQA and KGQA is the implementation of reasoning path (Refer to Section 3). Detailed prompts of all modules is in Appendix G.
@@ 9             Go to Line 11
Reasoning path for TableQA starts from a given table, to some columns and then to some rows, in order to constrain specific items in the table to answer the question. A sample of WTQ Pasupat and Liang (2015) dataset is shown in Table 6. The reasoning path of such question is from this table to column ‘‘Name in English’’ and ‘‘Depth’’, and then to row items whose ‘‘Name in English’’ is ‘‘Lake Tuz’’ and ‘‘Lake Palas Tuzla’’.
For reasoning path generation, we ask LLMs to generate a reasoning path, given a question and some table descriptions (i.e., the header and a randomly-sampled row), by in-context learning. Specifically, we require the LLMs to pick up at least to rows for complex arithmetic and aggregation operations in TableQA. For the example in Table 6, ideally, the LLM would generate a dict indicating the chosen header is “[’Name in English’, ’Depth’]”, and the row constrain is “"Name in English": [’Lake Tuz’, ’Lake Palas Tuzla’]”.
For reasoning path instantiation, since schemas of a table is not as massive as a KG, we don’t need a relation binding module to retrieve candidate relations and connect them according to the reasoning path. Instead, we first filter out columns and then filter out the rows in the reasoning path. For the example in Table 6, we filter out the column "Name in English" and "Depth" and rows whose "Name in English" is "Lake Tuz" or "Lake Palas Tuzla". If the the instantiation goes wrong, this is the necessary time to invoke editing. Here we only consider whether the instantiation of columns in the reasoning path goes wrong. If a selected row in the reasoning path fails to be instantiated, we do not invoke editing and return all rows in the columns.
For reasoning path editing, we collect the currently instantiated columns and candidate columns when the instantiation goes wrong, as error messages. Note that we only consider wrong columns as error, where we categorize the reasons of error as follows: i. “irrelevant column $\boldsymbol{col_{err}}$”: the column $\boldsymbol{col_{err}}$ in the reasoning path failed to be instantiated in the given table; ii. “insufficient columns in reasoning path”: the output reasoning path contains less than two columns (most questions in WTQ Pasupat and Liang (2015) and WikiSQL Zhong et al. (2017) need at least two columns to constrain the answer). Then, we ask the LLMs to edit previous reasoning path according to the feedback. For example, if chosen header is ["English Name"], not matching the headers in Table 6. We provide all candidate headers and a randomly-sampled row in Table 6 as error messages.
The QA reasoning is similar to KGQA, we concatenate the question and instantiated table items and ask the LLMs to answer the question.
@@ B.1 Reasoning Path Instantiation Details
In this part, we demonstrate more details of path-connecting mentioned in Section 4.3. Assume that we have a reasoning path starting from an entity, with $m$ relations $R$. All relations are bound to candidates $\hat{R}$. For the example in Figure 3, we have a reasoning path ‘‘[Nijmegen] serve_airport $\rightarrow$ contain’’ and have ‘‘$r_{1}:$ serve_airport’’ bound to “$\hat{r}_{1}:$ airport, terminal, serving port” in KG, and similarly have ‘‘$r_{2}:\leavevmode\nobreak$contain’’ bound to “$\hat{r}_{2}:\leavevmode\nobreak$contained by, contains, place of”. The purpose if path-connecting is to check if there exists an instance in KG where the starting point is entity ‘‘Nijmegen’’, then connect to some relations in $\hat{r}_{1}$, and then connect to some relations in $\hat{r}_{2}$.
The path-connecting is essentially an Breadth-first search (BFS) algorithm, with a queue $que$ containing possible entities connected to current relations. Firstly, we push the starting point (e.g., ‘‘Nijmegen’’) into $que$. Second, at each time, we pop out current head of $que$ and check if current candidate relations (e.g.,adjoin, terminal, serving port”) can connect to it on KG. If so, we successfully instantiate ‘‘serve_airport’’ on KG, and we can obtain some entities which are ‘‘serve_airport’’ of ‘‘Nijmegen’’ (e.g., ‘‘WZ air.’’ and ‘‘NTA.’’). Then, we add these entities to $que$. Such BFS searching continues until all relations are instantiated or anything goes wrong.